OpenStudy (anonymous):

Photoelectrons with a maximum speed of 7.00 · 105 m/s are ejected from a surface in the presence of light with a frequency of 8.00 · 1014Hz. Calculate the energy required to eject photoelectrons from the surface.

OpenStudy (anonymous):

@Nali

OpenStudy (nali):

do you mean the speed is 7 X 10^5 and frequency is 8 X 10^14 ?

OpenStudy (anonymous):

Yes, sorry, it didn't transfer over quite right.

OpenStudy (nali):

ok so 1st step is to find the kinetic energy KE= 1/2 mass x velocity ^2 = 1/2 mv^2 so you plug in the given velocity with the mass of the electron which is 9.11x10-31kg

OpenStudy (nali):

2nd step is to plug in the kinetic energy in the equation of threshold frequency hf= Work + KE So h is a constant that equals 6.626 x 10^-34 J-s and the frequency is given so you have to find work of ejecting photoelectrons.

OpenStudy (nali):

Work is going to be in Joules (J) and that would be your answer 

OpenStudy (nali):

@historygeek96 when you get ur final answer just tell me to make sure

OpenStudy (nali):

The mass of an electron is supposed to be 9.11x10^-31kg sorry, I forget to put the carrot there

OpenStudy (anonymous):

Thanks, I found the same thing in my lesson

OpenStudy (nali):

ok great!

OpenStudy (anonymous):

-2.2295x10^-19

OpenStudy (anonymous):

Unfortunately my options are 3.1x10^-19 and 7.5x10^19

OpenStudy (nali):

ok let me do it, just give me a moment

OpenStudy (anonymous):

Okay thank you so much

OpenStudy (nali):

ok I think you stopped after finding the KE and you did not continue by doing hf-KE to find the work in ejecting the photoelectrons

OpenStudy (nali):

KE=2.23 X 10^-19 but why do you have a negative sign there?

OpenStudy (anonymous):

hf is planck's constant right?

OpenStudy (nali):

yes

OpenStudy (anonymous):

I did continue past finding KE

OpenStudy (anonymous):

I keep redoing it and I keep getting the same thing...

OpenStudy (anonymous):

Oh I just got 3.074500e-19

OpenStudy (anonymous):

Which is one of the options I'm given

OpenStudy (nali):

I know wht you did your calculator is in radians mode

OpenStudy (anonymous):

Actually, I wasn't using frequency, I thought hf stood for planck's constant, not planck's constant multiplied by the frequency

OpenStudy (anonymous):

I thought it was the calculator too but I used two of them and got the same answers.

OpenStudy (nali):

oh ok so I know the answer now, what do you have?

OpenStudy (anonymous):

It was 3.1x10^-19.

OpenStudy (nali):

yes good job that is right

OpenStudy (anonymous):

Yeah, its an online class so once I've turned it in its graded instantly.